OpenStudy (anonymous):

need help with this equation q+3/3 + 9-5/5= 8/3

OpenStudy (anonymous):

q+1+9-1 = 8/3 q+9 = 8/3 q = 8/3 - 27/3 q = -19/3

OpenStudy (anonymous):

thank u